OpenStudy (anonymous):

what is the inverse of f(x)=3/x+5

OpenStudy (zehanz):

First, it must be clear whether you mean:\[f(x)=\frac{ 3 }{ x+5 }\]or:\[f(x)=\frac{ 3 }{ x }+5\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

f(x)=3/x+5

OpenStudy (zehanz):

A quick and dirty method is to use y instead of x:\[y=\frac{ 3 }{ x }+5\]Then switch x and y:\[x=\frac{ 3 }{ y }+5\]and then solve for y (rewrite this formula to read y=... again)

OpenStudy (anonymous):

ll be interchange the variable e(x) put x and x put f(x)=y

OpenStudy (anonymous):

ll be

OpenStudy (zehanz):

@mustry: that was what I wrote...

OpenStudy (anonymous):

ll be\[x=3/y+5\]

OpenStudy (zehanz):

I think @alanis should get the chance to try it now

OpenStudy (anonymous):

ur r8@ZeHanz

OpenStudy (anonymous):

wouldnt be